What is the angular acceleration of a ball that starts at rest and increases its angular velocity uniformly to rad/s in 10 s?

1
Expert's answer
2021-12-06T09:43:34-0500
"\alpha=\frac{1}{10}=0.1\frac{rad}{s^2}"